OpenStudy (unklerhaukus):

Continuous probability distribution over an angle range. (Q. in comments)

OpenStudy (unklerhaukus):

find the probability density of θ \[0<\theta<\pi , \rho(\theta)=k\] \[\int\limits_{0}^{π}\rho(\theta)d \theta=1 \] sum of probabilities is 1 \[ \int\limits_{0}^{π}d \theta=1/k\] \[\theta]π,0 = 1/k\] \[π=1/k\] \[k=1/π\] :.\[\rho(\theta)=(1/π), 0<\theta<\pi\]

OpenStudy (unklerhaukus):

the question i have is how would i go about finding the probability distribution for the projection onto the x-axis \[\rho(x)= ?\] dx=?

OpenStudy (unklerhaukus):

x=rcos(θ) dx=-rsin(θ)dθ